# Inferring – Evolution from the Gravitational Wave Background

### v1

@@ 3.4 Influence of the GSMF
In Figure 6 we plot against the – amplitude at a fixed redshift , locally all values of are equivalent and the value of we choose only shifts the evolving – models left or right on the x-axis. The color differentiates evolving (red) models from non-evolving (blue) and the marker styles show the pairs of models with the same prior set up aside from – evolution (i.e., each marker style appears twice, one on red and one on blue). The clear negative trend demonstrates the degeneracy between an increased number density of galaxies and an increased SMBH–galaxy mass ratio in our posterior models. In the cases where the GSMF was allowed to vary, the evolving models (Le11ev, Le03ev, Le00ev, and LM11ev) have a lower value for and a higher value for than their non-evolving counterparts (Le11ne and Le03ne). These models follow the same trend, but are additionally boosted relative to the fiducial J. Leja et al. (2020) GSMF by design (as described in section 2.3 and Appendix A). The only outlier to this trend is Le00ne. The best-fit likelihood value for this model is several orders of magnitude smaller than for other models (represented by the size of the points) and therefore its location on the plot is not an accurate reflection of the overall trend. Therefore we conclude that the high number density of massive SMBHs, required to reproduce the GWB, can be produced either by a large increase to galaxy number density (GSMF), a large increase to the SMBH–galaxy mass ratio, or a moderate increase to both.
@@ Appendix A GSMF Evolution Parameters and Scaling Relation Choice
In Figure 10 we demonstrate the impact of scaling relation choice on the GWB spectrum. The version of the scaling relation from J. Kormendy & L. C. Ho (2013) generally produces bigger SMBHs (and therefore higher GWB amplitudes) than other options (see, e.g., J. Simon, 2023). E. R. Liepold & C.-P. Ma (2024) use the N. J. McConnell & C.-P. Ma (2013) version of the – relation, which generally produces lower GWB amplitudes compared to J. Kormendy & L. C. Ho (2013). We additionally show the GWB spectrum using the scaling relation from F. Shankar et al. (2016).
This comparison demonstrates that, wherever our spectra fall short of the GWB data, using a relation with a lower amplitude would only exacerbate this problem. This also has implications for our choices for the evolution we approximate for the E. R. Liepold & C.-P. Ma (2024) GSMF. The moderately evolving GSMF would produce a much lower GWB amplitude if we use the N. J. McConnell & C.-P. Ma (2013) or F. Shankar et al. (2016) relations and so either the evolution would need to be more similar to that of the weakly evolving GSMF or a higher-amplitude – relation is needed.
Appendix B Corner Plots
@@ Appendix C Reliability of Methods
In this work we discovered a minor limitation of the direct-likelihood method which impacted the Le00ne model. It is more difficult for the posterior distributions to deviate from the priors because extreme values (e.g. more than 3 above or below the prior median) have a low chance of being randomly sampled, regardless of how well that value may describe the data. We see this limitation in our model with the lowest number of sampled parameters, Le00ne. The best-fit GWB spectrum from this model returned a notably low log-likelihood compared to the other models and produces a worse overall fit (see Figures 3 and 6). We can see, from Le03ev and Le00ev, that a value of is needed to reproduce the GWB amplitude without increasing galaxy number density. If the – relation is not allowed to evolve, the prior for would need to be sampled at a value away from the prior mean (assuming a normal distribution where ), and therefore extremely unlikely / impossible to be sampled. Without interpolating, any value that is not sampled simply does not have a corresponding GWB spectrum and the best-fit parameters may therefore not describe the data. This effect can be mitigated by choice of priors, but prior choice has its own impact on models and so future studies should consider the impact of this effect when using this method.

### v2

@@ 3.4 Influence of the GSMF
In Figure 6 we plot against the – amplitude at a fixed redshift , locally all values of are equivalent and the value of we choose only shifts the evolving – models left or right on the x-axis. The color differentiates evolving (red) models from non-evolving (blue) and the marker styles show the pairs of models with the same prior set up aside from – evolution (i.e., each marker style appears twice, one on red and one on blue). The clear negative trend demonstrates the degeneracy between an increased number density of galaxies and an increased SMBH–galaxy mass ratio in our posterior models. In the cases where the GSMF was allowed to vary, the evolving models (Le11ev, Le03ev, Le00ev, and LM11ev) have a lower value for and a higher value for than their non-evolving counterparts (Le11ne and Le03ne). These models follow the same trend, but are additionally boosted relative to the fiducial J. Leja et al. (2020) GSMF by design (as described in section 2.3 and Appendix A). The only outlier to this trend is Le00ne. The best-fit likelihood value for this model is several orders of magnitude smaller than for other models (represented by the size of the points) indicating it is a poor fit to the GWB data. While in good agreement with the GSMF (low ) its location on the plot is not an accurate reflection of the overall trend since the trend is defined by models that can reproduce the GWB. Therefore we conclude that the high number density of massive SMBHs, required to reproduce the GWB, can be produced either by a large increase to galaxy number density (GSMF), a large increase to the SMBH–galaxy mass ratio, or a moderate increase to both.
@@ Appendix A GSMF Evolution Parameters and Scaling Relation Choice
Appendix B Choice of Priors and Prior Tests
We chose astrophysically motivated priors as inputs to all our models. We investigated the impact that prior choice has on the models and found that our results are not sensitive to prior choice.
Several of our parameters, e.g., galaxy merger timescales and SMBH binary hardening timescales, do not have strong observational constraints. Instead, we make use of theoretically motivated prescriptions. The two studies we base our priors on both make use of the Illustris cosmological hydrodynamical simulations (M. Vogelsberger et al., 2014). We use the galaxy merger rates due to V. Rodriguez-Gomez et al. (2015) which provides an analytical prescription for merger rates that depends on galaxy mass and redshift. This is a different model to that used in G. Agazie et al. (2023b) though we find that the impact on the resulting GWB spectrum is not significant. We ran a series of tests allowing only the galaxy merger rate parameters to vary and found that the posteriors always recovered the priors (and therefore the “best fit” model was identical to the prior model). We concluded that our models have minimal to no constraining power over these parameters hence they are fixed for Le03ne, Le03ev, Le00ne, and Le00ev.
Numerous studies have investigated evolution of the physical separation of SMBH binaries, i.e., the “hardening” rate (M. C. Begelman et al., 1980; G. D. Quinlan, 1996; G. D. Quinlan & L. Hernquist, 1997; Q. Yu, 2002; A. Sesana & F. M. Khan, 2015; L. Z. Kelley et al., 2017; A. K. Bhowmick et al., 2024; S. Buttigieg et al., 2025, Harris et al. in prep). SMBH binary hardening rates are difficult to constrain observationally and so these studies rely on hydrodynamic and/or N-body simulations. Modeling the entire SMBH merger process is computationally challenging, however, leaving large uncertainties in hardening models (L. Z. Kelley et al., 2017). For this work, we use uniform priors on all free hardening parameters in our models. We see that the posterior distributions often push up against the limits of our prior range. In our testing, we found that the qualitative results of this work were not sensitive to the prior ranges we assume for these parameters. For example the hardening timescale, lower limit is 0 Gyr and cannot go lower despite the posterior distribution pushing against this lower limit. Additionally the parameter lower limit is roughly at the limit at which the timescale of the stellar scattering regime is minimized and lower values of have a negligible impact on the GWB spectrum. Finally the third hardening parameter we sample is the characteristic radius, , which indicates the point at which the dominant hardening mechanism transitions from dynamical friction to stellar scattering. This radius corresponds to the sphere of influence of the SMBHs and is on the order of 10s of pc. Our priors range from 2–20 pc, which is physically motivated (L. Z. Kelley et al., 2017). Allowing for larger values of can be useful for studies focused on constraints at small binary separations, but these parameters are not the focus of this study so we chose a realistic, but uniform range of values. We also note that there is a degeneracy between and such that large values of correspond to larger values of . Therefore changing the prior range on may affect the posterior values for, e.g., , but it does not notably impact the results of GSMF or – parameters. Improving the hardening parameter prior choice for GWB studies requires a dedicated multimessenger study. This is outside the scope of this work, but will be included in future investigations (e.g., Harris et al. in prep).
For the GSMF and – relation, we used models based on electromagnetic observations (J. Leja et al., 2020; J. Kormendy & L. C. Ho, 2013; E. R. Liepold & C.-P. Ma, 2024). To evaluate the constraining power of our models over the GSMF, we ran a model sampling only the 11 GSMF parameters with uniform priors (Le11Une) and repeated this model with a free parameter (Le11Uev). We found that the results of this test were consistent with the results of our study. The posterior distributions deviated from the uniform priors and produced BHMFs that are consistent with each other and also the BHMFs from our models with informed priors. The BHMFs from these models are shown in Figure 11, we additionally over-plot the median BHMF from the Le_ne models from Figure 7. We see that all three BHMFs return remarkably similar number densities in all redshift bins though Le11Une falls slightly short of the two evolving – models, a pattern also present in Figure 7. The posterior – evolution parameter from Le11Uev is and the distribution is 82.8% positive. While this is not statistically significant evolution, we believe this result supports our overall conclusions and is a good indicator that our models are not prior dominated for the GSMF parameters.
@@ Appendix B Choice of Priors and Prior Tests
Similar to how we tested different options for the GSMF, we additionally considered the effect of the version of the – relation we use. In Figure 10 we demonstrate the impact of scaling relation choice on the GWB spectrum. The version of the scaling relation from J. Kormendy & L. C. Ho (2013) generally produces bigger SMBHs (and therefore higher GWB amplitudes) than other options (see, e.g., J. Simon, 2023). E. R. Liepold & C.-P. Ma (2024) use the N. J. McConnell & C.-P. Ma (2013) version of the – relation, which generally produces lower GWB amplitudes compared to J. Kormendy & L. C. Ho (2013). We additionally show the GWB spectrum using the scaling relation from F. Shankar et al. (2016). This comparison demonstrates that, wherever our spectra fall short of the GWB data, using a relation with a lower amplitude would only exacerbate this problem. This also has implications for our choices for the evolution we approximate for the E. R. Liepold & C.-P. Ma (2024) GSMF.
@@ Appendix C Reliability of Methods
In this work we discovered a minor limitation of the direct-likelihood method which impacted the Le00ne model. It is more difficult for the posterior distributions to deviate from the priors because extreme values (e.g. more than 3 above or below the prior median) have a low chance of being randomly sampled, regardless of how well that value may describe the data. We see this limitation in our model with the lowest number of sampled parameters, Le00ne. The best-fit GWB spectrum from this model returned a notably low log-likelihood compared to the other models which means it produces a worse overall fit (see Figures 3 and 6). We can see, from Le03ev and Le00ev, that a value of is needed to reproduce the GWB amplitude without increasing galaxy number density. If the – relation is not allowed to evolve, the prior for would need to be sampled at a value away from the prior mean (assuming a normal distribution where ), and therefore extremely unlikely / impossible to be sampled. Without interpolating, any value that is not sampled simply does not have a corresponding GWB spectrum and the best-fit parameters may therefore not describe the data. This effect can be mitigated by choice of priors, but prior choice has its own impact on models and so future studies should consider the impact of this effect when using this method.
It is important to emphasize that the likelihoods we calculate are a measure of how well a given best-fit spectrum describes the observed GWB spectrum, but they are not marginalized over the respective parameter spaces. This means that the likelihoods to not contain information about prior-posterior agreement/disagreement and therefore do not indicate goodness of fit to EM data. These likelihoods cannot be used alone to evaluate which models are better models of both the GWB and EM data. To evaluate consistency with EM data we make use of other quantities such as the Kullback-Leilbler divergence, , and GSMF boost, . The “best” models have high likelihoods (good fits to the GWB) in addition to low and values (good agreement with the EM data).
Appendix D Corner Plots
Fig. Set1. Model Output Corner Plots